# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ Appendix D Proof of Lemma 10
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ be pure states. We define $\psi_{g}\coloneqq\mathcal{U}_{g}(\psi)$ and $\phi_{g}\coloneqq\mathcal{U}^{\prime}_{g}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi_{g}}$ and $\mathcal{Q}^{\phi_{g}}$. If $r>0$ satisfies $\mathcal{Q}^{\psi_{g}}\geq r\mathcal{Q}^{\phi_{g}}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
@@
Following the standard argument using typical sequence [72, 21], we prove Eq. (75), i.e., $\sum_{i\in[k]}p_{i}\mathcal{A}_{\mathrm{c}}(\rho_{i})\geq\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho=\sum_{i\in[k]}p_{i}\rho_{i}$, where $[k]\coloneqq\{1,2,\cdots,k\}$.
@@
$\leq\max_{\bm{i}\in\mathcal{T}_{\delta}}T\left(\mathcal{E}_{N}^{(\bm{i})}(\phi^{\otimes\lceil rN\rceil}),\rho_{\bm{i}}\right)$ (366)
@@
Appendix F Convex roof of QGT
@@
$\mathcal{Q}^{\ket{0}}=\mathcal{Q}^{\ket{1}}=\begin{pmatrix}1&\mathrm{i}&0\\ -\mathrm{i}&1&0\\ 0&0&0\end{pmatrix},\quad\mathcal{Q}^{\psi_{\pm}}=\begin{pmatrix}\cos^{2}\varphi&\mathrm{i}\cos\varphi&\mp\cos\varphi\\ -\mathrm{i}\cos\varphi&1&\pm\mathrm{i}\sin\varphi\\ \pm\cos\varphi\sin\varphi&\mp\mathrm{i}\sin\varphi&\sin^{2}\varphi\end{pmatrix}.$ (369)
For the first decomposition $\rho=\epsilon\ket{0}\bra{0}+(1-\epsilon)\frac{1}{2}I$, the average QGT is given by
$\epsilon\mathcal{Q}^{\ket{0}}=\begin{pmatrix}\cos\varphi&\mathrm{i}\cos\varphi&0\\ -\mathrm{i}\cos\varphi&\cos\varphi&0\\ 0&0&0\end{pmatrix},$ (370)
while the average QGT for the second decomposition $\rho=\frac{1}{2}(\ket{\psi_{+}}\bra{\psi_{+}}+\ket{\psi_{+}}\bra{\psi_{+}})$ is given by
@@
Appendix G QGT of purification
For a mixed state $\rho$ on a system of interest $S$, we consider minimizing the QGT over the set of all purifications |$\ket{\Psi_{\rho}}$ of $\rho$, by extending the arguments in [21]. Let $A$ denote the ancillary system added to purify $\rho$. For sets of Hermitian operator $\bm{X}_{S}$ and $\bm{X}_{A}$, we define
@@ Appendix G QGT of purification
For two different purifications |$\ket{\Phi_{\rho}}$ and |$\ket{\Phi_{\rho}^{\prime}}$ of $\rho$, there exists a unitary operator $U_{A}$ on $A$ such that |$\ket{\Phi_{\rho}^{\prime}}=I_{S}\otimes U_{A}\ket{\Phi_{\rho}}$. Since
$\braket{\Phi_{\rho}^{\prime}}{O_{tot}(I-\Phi_{\rho})O_{tot}^{\dagger}}{\Phi_{\rho}^{\prime}}=\braket{\Phi_{\rho}}{O_{tot}^{\prime}(I-\Phi_{\rho})O_{tot}^{\prime{\dagger}}}{\Phi_{\rho}},$ (374)
where $O_{tot}^{\prime}\coloneqq O_{S}\otimes I_{A}+I_{S}\otimes\bm{\gamma}^{\dagger}\bm{X}_{A}^{\prime}$ and $X_{A,\mu}^{\prime}\coloneqq U_{A}^{\dagger}X_{A,\mu}U_{A}$, it suffices to minimize over possible $\bm{X}_{A}$ for a fixed purification. In the following, we denote by |$\ket{\Phi_{\rho}}$ a purification of $\rho$ given by
[…]
where $\rho=\sum_{i}p_{i}\ket{\phi_{i}}\bra{\phi_{i}}$ | is the eigenvalue decomposition of $\rho$.
By redefining
$X_{A,\mu}\to\tilde{X}_{A,\mu}\coloneqq X_{A,\mu}-\braket{\Phi_{\rho}}{X_{S,\mu}\otimes I_{A}+I_{S}\otimes X_{A,\mu}}{\Phi_{\rho}}I_{A},$ (376)
we get
$\braket{\Phi_{\rho}}{X_{S,\mu}\otimes I_{A}+I_{S}\otimes\tilde{X}_{A,\mu}}{\Phi_{\rho}}=0$ (377)
@@
$\delta V\left(\Phi_{\rho},O_{tot}\right)$ $=\sum_{i}p_{i}\braket{\phi_{i}}{O_{A}(\delta O_{A})^{\dagger}}{\phi_{i}}+\sum_{i}p_{i}\braket{\phi_{i}}{(\delta O_{A})O_{A}^{\dagger}}{\phi_{i}}$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}}{(O_{S}\otimes(\delta O_{A})^{\dagger}+O_{S}^{\dagger}\otimes(\delta O_{A}))}{\phi_{j}\otimes\phi_{j}}+O((\delta O_{A})^{2})$ (379)
@@
$\mathcal{K}_{\mu\nu}$ $\coloneqq\sum_{i}p_{i}\braket{\phi_{i}}{X_{A,\mu}(\delta X_{A,\nu})}{\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}}{X_{S,\mu}\otimes(\delta X_{A,\nu})}{\phi_{j}\otimes\phi_{j}}$
$\quad+\sum_{i}p_{i}\braket{\phi_{i}}{(\delta X_{A,\mu})X_{A,\nu}}{\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}}{X_{S,\nu}\otimes(\delta X_{A,\mu})}{\phi_{j}\otimes\phi_{j}}.$ (381)
@@
$K_{\mu\mu}=0$ holds for all $\delta\bm{X}_{A}$ only if
[…]
@@
$X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}}{X_{S,\mu}}{\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}.$ | . (383)
@@
$\mathcal{M}_{2,1}$ $=\sum_{i}p_{i}\braket{\phi_{i}}{X_{A,2}(\delta X_{A,1})}{\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}}{X_{S,2}\otimes(\delta X_{A,1})}{\phi_{j}\otimes\phi_{j}},$ (384)
which does not vanish for $X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}}{X_{S,\mu}}{\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}$ |. Therefore, for $\dim G>1$, the minimum of QGT of purification does not exist in general.
@@
$=\sum_{i}\left(\frac{1}{1-q}p_{i}\braket{\phi_{i}}{O_{A}(\delta O_{A})^{\dagger}}{\phi_{i}}+\frac{1}{q}p_{i}\braket{\phi_{i}}{(\delta O_{A})^{\dagger}O_{A}}{\phi_{i}}\right)$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\left(\frac{1}{1-q}+\frac{1}{q}\right)\braket{\phi_{i}\otimes\phi_{i}}{O_{S}\otimes(\delta O_{A})^{\dagger}}{\phi_{j}\otimes\phi_{j}}$
$\quad+\sum_{i}\left(\frac{1}{1-q}p_{i}\braket{\phi_{i}}{(\delta O_{A})O_{A}^{\dagger}}{\phi_{i}}+\frac{1}{q}p_{i}\braket{\phi_{i}}{O_{A}^{\dagger}(\delta O_{A})}{\phi_{i}}\right)$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\left(\frac{1}{1-q}+\frac{1}{q}\right)\braket{\phi_{i}\otimes\phi_{i}}{O_{S}^{\dagger}\otimes(\delta O_{A})}{\phi_{j}\otimes\phi_{j}}+O((\delta O_{A})^{2})$ (387)

### v3

@@ Appendix D Proof of Lemma 10
[…]
@@
Following the standard argument using typical sequence [91, 24], we prove Eq. (79), i.e., $\sum_{i\in[k]}p_{i}\mathcal{A}_{\mathrm{c}}(\rho_{i})\geq\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho=\sum_{i\in[k]}p_{i}\rho_{i}$, where $k]\coloneqq\{1,2,\cdots,k\}$.
@@
$\leq\max_{\bm{i}\in\mathcal{T}_{\delta}}T\left(\mathcal{E}_{N}^{(\bm{i})}(\phi^{\otimes\lceil rN\rceil}),\rho_{\bm{i}}\right),$ (410)
@@
Appendix F Proof of an upper bound on the asymmetry cost
In this section, we present the proof of Proposition 14, which provides an upper bound on the asymmetry cost. It is worth emphasizing that, unlike the result in the previous section, Proposition 14 does not impose any constraint on the symmetry subgroups in the ensemble decomposition. Intuitively, this relaxation is possible because the symmetry subgroup constraint is only relevant for the estimation step in the estimation-and-conversion strategy introduced in Sec.V, while the conversion step succeeds as long as the monotonicity condition of QGTs is satisfied. Below, we carefully modify the argument of typical sequences presented in the previous section, and use it to extend the proof of Proposition 6, thereby establishing Proposition 14. Following the convention in the main text, we define $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ for a unitary representation $U$ of $G$. Note that, with a slight abuse of notation, we also write $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ for $\bm{\theta}\in\mathbb{R}^{\dim G}$.
Lemma 42.
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\rho$ be a state on $\mathcal{H}$ and $\rho$ be a state on $\mathcal{H}^{\prime}$. Consider an ensemble $\{p_{i},\sigma_{i}\}_{i\in[k]}$ such that $\sigma=\sum_{i}p_{i}\sigma_{i}$, where $k]\coloneqq\{1,\cdots,k\}$. Suppose that for each $i$, for some finite $r_{i}>0$, there exists a sequence of quantum channels $\mathcal{E}_{N}^{(i,g)}$ such that
$\delta_{N}^{(i)}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(i,g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil r_{i}N\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{i,g}\right)^{\otimes N}\right)$ (413)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}^{(i)}(g,\bm{u})=0$ for $\epsilon^{\prime}\in(0,\epsilon)$, where $\rho_{g}\coloneqq\mathcal{U}_{g}(\rho)$ and $\sigma_{i,g}\coloneqq\mathcal{U}_{g}^{\prime}(\sigma_{i})$. Then there exists a sequence of quantum channels $\mathcal{E}_{N}^{(g)}$ and $\epsilon^{\prime}\in(0,1)$ such that
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)$ (414)
for $r\coloneqq\sum_{i}p_{i}r_{i}+\Delta$ and $\sigma_{g}\coloneqq\mathcal{U}_{g}^{\prime}(\sigma)$, where $\Delta$ is an arbitrary positive number.
Proof.
If $p_{i}=1$ for some $i$, then the statement is trivial. In the following, we only consider the case where $p_{i}<1$ for all $i$. We take $\delta>0$ such that $p_{i}+\delta<1$ for all $i$.
The i.i.d. copies of $\sigma$ are expressed as
$\sigma^{\otimes N}=\left(\sum_{i\in[k]}p_{i}\sigma_{i}\right)^{\otimes N}=\sum_{\bm{i}\in[k]^{N}}p_{\bm{i}}\rho_{\bm{i}},$ (415)
where $\bm{i}\coloneqq(i_{1},\cdots,i_{N})\in[k]^{N}$, $p_{\bm{i}}\coloneqq\prod_{j=1}^{N}p_{i_{j}}$, and $\sigma_{\bm{i}}\coloneqq\otimes_{j=1}^{N}\rho_{i_{j}}$. For $l\in[k]$, let $n_{l}(\bm{i})$ denote the number of occurrence of $l$ in $(i_{1},\cdots,i_{N})$. For $\delta>0$, we define $\delta$-typical sequence as
$\mathcal{T}_{\delta}\coloneqq\left\{\bm{i}=(i_{1},i_{2},\cdots,i_{N})\in[k]\,\middle|\,\forall l\in[k],\,\left|\frac{n_{l}(\bm{i})}{N}-p_{l}\right|\leq\delta\right\}.$ (416)
Introducing
$\tilde{\sigma}_{N}\coloneqq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\sigma_{\bm{i}}+p_{\mathrm{err}}\sigma_{\mathrm{sym}},$ (417)
where $p_{\mathrm{err}}\coloneqq 1-\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}$ and $\sigma_{\mathrm{sym}}$ is some symmetric state (,e.g., the maximally mixed state), the standard argument on a typical sequence implies
$\lim_{N\to\infty}T\left(\sigma^{\otimes N},\tilde{\sigma}_{N}\right)=0.$ (418)
From the invariance of the trace distance under unitary operations, it also implies
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{g})\right)^{\otimes N},\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime\otimes N}\circ\mathcal{U}_{g}^{\prime\otimes N}\right)\left(\tilde{\sigma}_{N}\right)\right)=0.$ (419)
Since the state $\sigma_{\bm{i}}$ is equivalent to $\bigotimes_{l\in[k]}\sigma_{l}^{\otimes n_{l}(\bm{i})}$ up to permutation, let us first consider a channel that approximately creates $\bigotimes_{l\in[k]}\sigma_{l}^{\otimes n_{l}(\bm{i})}$ for a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$. Define a sequence of quantum channels $\mathcal{F}_{N}$ by
$\mathcal{F}_{N}^{(g)}\coloneqq\bigotimes_{l\in[k]}\left(\mathcal{E}^{(l,g)}_{N_{l}}\right)$, (420)
where we have introduced $N_{l}\coloneqq\lceil N(p_{l}+\delta)\rceil$. From the subadditivity of the trace distance, we get
$T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}T\left(\mathcal{E}^{(l,g)}_{N_{l}}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})^{\otimes N_{l}}\right)$
$=\sum_{l\in[k]}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\frac{\sqrt{N_{l}}}{\sqrt{N}}\right).$ (421)
Since $\delta$ satisfies $p_{l}+\delta<1$, we have $N_{l}\leq N$, implying that $\frac{\sqrt{N_{l}}}{\sqrt{N}}\leq 1$. For all sufficiently large $N$, $N^{\epsilon^{\prime}}<N_{l}^{\epsilon}$ holds for $\epsilon^{\prime}\in(0,\epsilon)$, meaning that
$\|\bm{u}\|<N^{\epsilon}\implies\left\|\bm{u}\frac{\sqrt{N_{l}}}{\sqrt{N}}\right\|<N_{l}^{\epsilon}.$ (422)
Therefore, we get
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes N_{l}}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right)$ (423)
holds for all sufficiently large $N$. Since $N_{l}\geq n_{l}(\bm{i})$ holds for a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$, we can introduce a channel $\mathcal{D}_{N_{l}\to n_{l}(\bm{i)}}$ that discards $N_{l}-n_{l}(\bm{i})$ subsystems if $N_{l}>n_{l}(\bm{i})$ and does nothing if $N_{l}=n_{l}(\bm{i})$. For $\mathcal{D}_{N}^{\bm{i}}\coloneqq\bigotimes_{l\in[k]}\mathcal{D}_{N_{l}\to n_{l}(\bm{i)}}$, we get
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{D}_{N}^{(\bm{i})}\circ\mathcal{F}_{N}^{(g)}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\sigma_{l,g})\right)^{\otimes n_{l}(\bm{i})}\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right).$ (424)
Denoting by $\mathcal{P}_{\bm{i}}$ a channel that swaps subsystems so that $\mathcal{P}^{(\bm{i})}(\bigotimes_{l\in[k]}\sigma_{l}^{\otimes n_{l}(\bm{i})})=\sigma_{\bm{i}}$, which commutes with $\mathcal{U}^{\otimes N}_{\bm{u}}\circ\mathcal{U}^{\otimes N}_{g}$, we get
$\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$
$\leq\sum_{l\in[k]}\sup_{g\in G}\sup_{\|\bm{u}\|<N_{l}^{\epsilon}}\delta_{N_{l}}^{(l)}\left(g,\bm{u}\right)$ (425)
for $\mathcal{E}^{(\bm{i},g)}_{N}\coloneqq\mathcal{P}^{(\bm{i})}\circ\mathcal{D}_{N}^{\bm{i}}\circ\mathcal{F}_{N}^{(g)}$.
We now fix $\Delta>0$ so that $\sum_{l}\lceil r_{l}N_{l}\rceil<\lceil rN\rceil$ for $r\coloneqq\sum_{l}p_{l}r_{l}+\Delta$ for all sufficiently large $N$. Since $\delta>0$ can be chosen arbitrarily small and $r_{l}$ is finite from the assumption, $\Delta$ can also be made arbitrarily small. Defining a sequence of channels $\{\mathcal{E}_{N}\}_{N}$ by
$\mathcal{E}_{N}^{(g)}(\cdot)\coloneqq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\mathcal{E}_{N}^{(\bm{i},g)}\circ\mathcal{D}_{\lceil rN\rceil\to\sum_{l}\lceil r_{l}N_{l}\rceil}(\cdot)+p_{\mathrm{err}}\mathcal{E}_{\mathrm{sym},N}(\cdot),$ (426)
where $\mathcal{E}_{\mathrm{sym},N}(\cdot)$ denotes a channel that maps an arbitrary state into a symmetric state $\rho_{\mathrm{sym}}$, we get
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil rN\rceil},\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime\otimes N}(\tilde{\sigma}_{N,g})\right)$ (427)
$=T\left(\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right)+p_{\mathrm{err}}\rho_{\mathrm{sym}},\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\prime}(\sigma_{\bm{i},g})\right)^{\otimes N}+p_{\mathrm{err}}\rho_{\mathrm{sym}}\right)$ (428)
$\leq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right)$ (429)
$\leq\max_{\bm{i}\in\mathcal{T}_{\delta}}T\left(\mathcal{E}^{(\bm{i},g)}_{N}\left(\bigotimes_{l\in[k]}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\rho_{g})\right)^{\otimes\lceil r_{l}N_{l}\rceil}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}^{\otimes N}(\sigma_{\bm{i},g})\right),$ (430)
where we have used the joint convexity of the trace distance and $\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\leq 1$. From Eqs. (419), (425), (430), we get
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\rho_{g}\right)^{\otimes\lceil rN\rceil}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\sigma_{g}\right)^{\otimes N}\right)=0.$ (431)
Appendix G Convex roof of QGT
@@
[…]
@@
Appendix H QGT of purification
[…]
@@ Appendix H QGT of purification
$V\left(\Phi_{\rho},O_{tot}\right)\coloneqq\braket{\Phi_{\rho}|O_{tot}(I-\Phi_{\rho})O_{tot}^{\dagger}|\Phi_{\rho}}=\braket{\Phi_{\rho}|O_{tot}O_{tot}^{\dagger}|\Phi_{\rho}}-\braket{\Phi_{\rho}|O_{tot}|\Phi_{\rho}}\braket{\Phi_{\rho}|O_{tot}^{\dagger}|\Phi_{\rho}}$ (436)
@@
$\braket{\Phi_{\rho}^{\prime}|O_{tot}(I-\Phi_{\rho})O_{tot}^{\dagger}|\Phi_{\rho}^{\prime}}=\braket{\Phi_{\rho}|O_{tot}^{\prime}(I-\Phi_{\rho})O_{tot}^{\prime{\dagger}}|\Phi_{\rho}},$ (437)
@@
where $|\rho=\sum_{i}p_{i}\ket{\phi_{i}}\bra{\phi_{i}}$ is the eigenvalue decomposition of $\rho$.
By redefining
$X_{A,\mu}\to\tilde{X}_{A,\mu}\coloneqq X_{A,\mu}-\braket{\Phi_{\rho}|X_{S,\mu}\otimes I_{A}+I_{S}\otimes X_{A,\mu}|\Phi_{\rho}}I_{A},$ (439)
we get
$\braket{\Phi_{\rho}|X_{S,\mu}\otimes I_{A}+I_{S}\otimes\tilde{X}_{A,\mu}|\Phi_{\rho}}=0$ (440)
for all $\mu$, implying that $\braket{\Phi_{\rho}|\bm{\gamma}^{\dagger}\bm{X}_{S}\otimes I_{A}+I_{S}\otimes\bm{\gamma}^{\dagger}\tilde{\bm{X}}_{A}|\Phi_{\rho}}=0$. Since adding the multiple of $I_{A}$ to $X_{A,\mu}$ does not change $V\left(\Phi_{\rho},O_{tot}\right)$, we assume, without loss of generality, that $\braket{\Phi_{\rho}|O_{tot}|\Phi_{\rho}}=\braket{\Phi_{\rho}|O_{tot}^{\dagger}|\Phi_{\rho}}=0$ in the following. In this case, we have
$V\left(\Phi_{\rho},O_{tot}\right)$ $=\sum_{i}p_{i}\braket{\phi_{i}|O_{S}O_{S}^{\dagger}|\phi_{i}}+\sum_{i}p_{i}\braket{\phi_{i}|O_{A}O_{A}^{\dagger}|\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}|(O_{S}\otimes O_{A}^{\dagger}+O_{S}^{\dagger}\otimes O_{A})|\phi_{j}\otimes\phi_{j}}.$ (441)
@@
$\delta V\left(\Phi_{\rho},O_{tot}\right)$ $=\sum_{i}p_{i}\braket{\phi_{i}|O_{A}(\delta O_{A})^{\dagger}|\phi_{i}}+\sum_{i}p_{i}\braket{\phi_{i}|(\delta O_{A})O_{A}^{\dagger}|\phi_{i}}$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}|(O_{S}\otimes(\delta O_{A})^{\dagger}+O_{S}^{\dagger}\otimes(\delta O_{A}))|\phi_{j}\otimes\phi_{j}}+O((\delta O_{A})^{2})$ (442)
@@
$\mathcal{K}_{\mu\nu}$ $\coloneqq\sum_{i}p_{i}\braket{\phi_{i}|X_{A,\mu}(\delta X_{A,\nu})|\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}|X_{S,\mu}\otimes(\delta X_{A,\nu})|\phi_{j}\otimes\phi_{j}}$
$\quad+\sum_{i}p_{i}\braket{\phi_{i}|(\delta X_{A,\mu})X_{A,\nu}|\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}|X_{S,\nu}\otimes(\delta X_{A,\mu})|\phi_{j}\otimes\phi_{j}}.$ (444)
@@
$\mathcal{K}_{\mu\mu}=0$ holds for all $\delta\bm{X}_{A}$ only if
$\braket{\phi_{i}|X_{A,\mu}|\phi_{j}}=-\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}|X_{S,\mu}|\phi_{i}},$ (445)
@@
$|.\displaystyle X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}|X_{S,\mu}|\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}$. (446)
@@
$\mathcal{K}_{2,1}$ $=\sum_{i}p_{i}\braket{\phi_{i}|X_{A,2}(\delta X_{A,1})|\phi_{i}}+\sum_{i,j}\sqrt{p_{i}p_{j}}\braket{\phi_{i}\otimes\phi_{i}|X_{S,2}\otimes(\delta X_{A,1})|\phi_{j}\otimes\phi_{j}},$ (447)
which does not vanish for $|X_{A,\mu}=-\sum_{i,j}\frac{2\sqrt{p_{i}p_{j}}}{p_{i}+p_{j}}\braket{\phi_{j}|X_{S,\mu}|\phi_{i}}\ket{\phi_{i}}\bra{\phi_{j}}$. Therefore, for $\dim G>1$, the minimum of QGT of purification does not exist in general.
@@
$=\sum_{i}\left(\frac{1}{1-q}p_{i}\braket{\phi_{i}|O_{A}(\delta O_{A})^{\dagger}|\phi_{i}}+\frac{1}{q}p_{i}\braket{\phi_{i}|(\delta O_{A})^{\dagger}O_{A}|\phi_{i}}\right)$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\left(\frac{1}{1-q}+\frac{1}{q}\right)\braket{\phi_{i}\otimes\phi_{i}|O_{S}\otimes(\delta O_{A})^{\dagger}|\phi_{j}\otimes\phi_{j}}$
$\quad+\sum_{i}\left(\frac{1}{1-q}p_{i}\braket{\phi_{i}|(\delta O_{A})O_{A}^{\dagger}|\phi_{i}}+\frac{1}{q}p_{i}\braket{\phi_{i}|O_{A}^{\dagger}(\delta O_{A})|\phi_{i}}\right)$
$\quad+\sum_{i,j}\sqrt{p_{i}p_{j}}\left(\frac{1}{1-q}+\frac{1}{q}\right)\braket{\phi_{i}\otimes\phi_{i}|O_{S}^{\dagger}\otimes(\delta O_{A})|\phi_{j}\otimes\phi_{j}}+O((\delta O_{A})^{2})$ (450)